# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v2

@@ Appendix E Proof of Lemma 3
The proof is similar to the proof of Theorem 3 in Zhang and Sugiyama [2023] and Lemma 1 of Lee and Oh [2024]. Define
$\widetilde{f}_{k,h}(\theta)=f_{k,h}(\theta_{k,h})+\langle g_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{H_{k,h}(\widetilde{\theta}_{k,h})},$
which is a second order approximation of the original function $f_{k,h}(\theta)$ at $\widetilde{\theta}_{k,h}$. Then, the update rule in (9) can be rewritten as
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{f}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}.$ (28)
Then, we present the following lemma that bounds the estimation error of the update rule in (28).
Lemma 12 (Lemma E.1 of Lee and Oh [2024]).
For the update rule defined in (28), it holds that
@@
$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}\leq$ $2\eta\left(\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})\right)+4\lambda B$
@@
We bound the right-hand side of the above lemma separately in the following.
First, we bound the first term. Define the softmax function as follows.
$[\sigma_{k,h}(z)]_{s}=\frac{\exp([z])_{s}}{1+\sum_{s\in\dot{\mathcal{S}}_{k,h}}\exp([z])_{s}},\quad\forall s\in\mathcal{S}_{k,h}.$
Then, the loss function in (2) can be rewritten as
@@
$f_{k,h}(z_{k,h},y_{k,h})=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\operatorname{\mathds{1}}[y_{k,h}^{s^{\prime}}=1]\log\left(\frac{1}{[\sigma_{k,h}(z_{k,h})]_{s^{\prime}}}\right).$
Define a pseudo-inverse function of $\sigma_{k,h}(\cdot)$ as
@@
$[\sigma_{k,h}^{-1}(p)]_{s^{\prime}}=\log\left(\frac{[p]_{s^{\prime}}}{1-\lVert p\rVert_{1}}\right),\forall p\in\{p\in[0,1]^{N_{k,h}}\mid\lVert p\rVert_{1}<1\}.$
Then, we decompose the first term as follows.
$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})$
=\ = $\underbrace{\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})}_{(a)}+\underbrace{\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})}_{(b)}$
@@
First, we bound the term $(a)$ as follows.
Lemma 13 (Lemma E.2 of Lee and Oh [2024]).
Let $\delta\in(0,1]$ and $\lambda\geq 1$, for all $k\in[K],h\in[H]$, with probability at least $1-\delta$, we have
$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})$
@@
Then, we bound the term $(b)$ as follows.
Lemma 14 (Lemma E.3 of Lee and Oh [2024]).
For any $c\geq 0$, let $\lambda\geq\max\{2,72cd\}$, then for all $k\in[K],h\in[H]$, we have
$\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})\leq\frac{1}{2c}\sum_{i=1}^{k}\left\|\widetilde{\theta}_{i,h}-\theta_{i+1,h}\right\|_{\mathcal{H}_{i,h}}^{2}+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right).$
Now, we are ready to prove Lemma 3. Combining Lemma 12, Lemma 13, and Lemma 14, we have

### v3

@@ D.1 Proof of Claim 1
First, by the definition of MNL mixture MDP, we have $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ and $s^{\prime}\in\mathcal{S}_{h,s,a}$, it holds that $p_{s,a}^{s^{\prime}}(\theta)\geq\exp(-B)/(U\exp(B)),\forall\theta\in\mathbb{R}^{d}$, thus $\kappa^{*}\geq\kappa\geq 1/(U\exp(2B))^{2}$. Next, consider the state-action pair $(s,a)$ at stage $h$ with the maximum number of reachable states $U$, it is clear that $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})=1$. This implies that $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})p_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})=1$. Applying the pigeonhole principle, there exists $s^{\prime},s^{\prime\prime}\in\mathcal{S}_{h,s,a}$ such that $p_{s,a}^{s^{\prime}}(\theta_{h}^{*})p_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\leq 1/U^{2}$. Thus, we conclude that $\kappa\leq\kappa^{*}\leq 1/U^{2}$. This finishes the proof. ∎
Appendix E Omitted Proofs for Section 4
E.1 Useful Lemma
Lemma 10.
[…]
Proof.
By the multivariate mean value theorem, we have
$\mathcal{G}_{k,h}(\theta_{1})-\mathcal{G}_{k,h}(\theta_{2})=\nabla\mathcal{L}_{k,h}(\theta_{1})-\nabla\mathcal{L}_{k,h}(\theta_{2})=\int_{0}^{1}\nabla^{2}\mathcal{L}_{k,h}(\theta_{2}+t(\theta_{1}-\theta_{2}))\mathrm{d}t(\theta_{1}-\theta_{2}).$
Hence, we have
$\lVert\mathcal{G}(\theta_{1})-\mathcal{G}(\theta_{2})\rVert_{G_{k,h}^{-1}(\theta_{1},\theta_{2})}=\lVert\theta_{1}-\theta_{2}\rVert_{G_{k,h}(\theta_{1},\theta_{2})}.$
where $G_{k,h}(\theta_{1},\theta_{2})=\int_{0}^{1}\nabla^{2}\mathcal{L}_{k,h}(\theta_{2}+t(\theta_{1}-\theta_{2}))\mathrm{d}t$. By self-concordant-like property of $\mathcal{L}_{k,h}$ in Proposition 1, we have $\mathcal{H}_{k,h}(\theta_{1})\preceq({1+3\sqrt{2}})G_{k,h}(\theta_{1},\theta_{2})$. As a result, we have
$\left\|\theta_{1}-\theta_{2}\right\|_{\mathcal{H}_{k,h}\left(\theta_{1}\right)}$ $\leq(1+3\sqrt{2})^{1/2}\left\|\theta_{1}-\theta_{2}\right\|_{G_{k,h}\left(\theta_{1},\theta_{2}\right)}$
$=(1+3\sqrt{2})^{1/2}\left\|\mathcal{G}_{k,h}\left(\theta_{1}\right)-\mathcal{G}_{k,h}\left(\theta_{2}\right)\right\|_{G_{k,h}^{-1}\left(\theta_{1},\theta_{2}\right)}$
$\leq(1+3\sqrt{2})\left\|\mathcal{G}_{k,h}\left(\theta_{1}\right)-\mathcal{G}_{k,h}\left(\theta_{2}\right)\right\|_{\mathcal{H}_{k,h}^{-1}\left(\theta_{1}\right)}.$
This finishes the proof. ∎
E.2 Proof of Lemma 1
Proof.
Since $\widehat{\theta}_{k,h}$ minimizes $\mathcal{L}_{k,h}(\theta)$, we have $\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})=\bm{0}$. Thus, we have
$\mathcal{G}_{k,h}(\theta_{h}^{*})-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})=\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta_{h}^{*})-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}+\lambda_{k}\theta_{h}^{*}.$
Therefore, since $\lVert\theta_{h}^{*}\rVert\leq B$ and $\mathcal{H}_{k,h}(\theta_{h}^{*})\succeq\lambda_{k}I$, for any $\delta\in(0,1]$, with probability at least $1-\frac{\delta}{H}$,
@@ E.2 Proof of Lemma 1
$\lVert\mathcal{G}_{k,h}(\theta_{h}^{*})-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}\leq$ $\Big{\lVert}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta_{h}^{*})-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}+\sqrt{\lambda_{k}}B$
[…]
where the last inequality holds by the Bernstein-type concentration inequality in Lemma 13. Then, by the determinant inequality in Lemma 14, we have $\det(\mathcal{H}_{k,h}(\theta_{h}^{*}))\leq\left(\lambda_{k}+\frac{k}{d}\right)^{d}$. Thus, we obtain
$\lVert\mathcal{G}_{k,h}(\theta_{h}^{*})-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}\leq\left(B+\frac{1}{2}\right)\sqrt{\lambda_{k}}+\frac{2}{\sqrt{\lambda_{k}}}\log\left(\frac{4}{\delta}\left(1+\frac{kH}{d\lambda_{k}}\right)\right).$
By the configuration that $\lambda_{k}=d\log(kH/\delta)$ and applying the union bound for $h\in[H]$, we have with probability at least $1-\delta$, for all $h\in[H]$ simultaneously, it holds that
$\lVert\mathcal{G}_{k,h}(\theta_{h}^{*})-\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}\leq(B+3)\sqrt{d\log(kH/\delta)}.$
This finishes the proof. ∎
E.3 Proof of Theorem 1
Proof.
By Lemma 10, with probability at least $1-\delta$, it holds that
$\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}$ $\leq(1+3\sqrt{2})\lVert\mathcal{G}_{k,h}(\widehat{\theta}_{k,h})-\mathcal{G}_{k,h}(\theta_{h}^{*})\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}\leq(1+3\sqrt{2})\widehat{\beta}_{k}.$
where the last inequality holds the confidence set $\widehat{\mathcal{C}}_{k,h}$ in Lemma 1. Then, by Lemma 7, we have
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{1st}}+\epsilon_{s,a}^{\mathtt{2nd}}.$
where
$\epsilon_{s,a}^{\mathtt{1st}}$ $=(1+3\sqrt{2})H\widehat{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{h}^{*})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})},$
$\epsilon_{s,a}^{\mathtt{2nd}}$ $=90H\widehat{\beta}_{k}^{2}\max_{s^{\prime}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}^{2}.$
By Lemma 9, we have
$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$ $\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}(\epsilon_{k,h}^{\mathtt{1st}}+\epsilon_{k,h}^{\mathtt{2nd}})+H\sqrt{2KH\log(2/\delta)}$ (20)
Next, we bound $\epsilon_{k,h}^{\mathtt{1st}}$ and $\epsilon_{k,h}^{\mathtt{2nd}}$ respectively.
Bounding $\epsilon_{k,h}^{\mathtt{1st}}$.  By statement (I) of Lemma 6, we have
$\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\theta_{h}^{*})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}$
[…]
By the configuration that $\widehat{\beta}_{k}=(B+3)\sqrt{d\log(k/\delta)}$, we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{1st}}\leq(1+3\sqrt{2})(B+3)dH^{2}\sqrt{K\log\left(1+\frac{K}{d\lambda_{K}}\right)\log\left(\frac{k}{\delta}\right)}.$ (21)
Bounding $\epsilon_{k,h}^{\mathtt{2nd}}$.  By statement (IV) of Lemma 6, we have
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{2nd}}$ $=90H^{2}\sum_{k=1}^{K}\widehat{\beta}_{k}^{2}\max_{s^{\prime}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}(\theta_{h}^{*})}^{2}$
$\leq\frac{180}{\kappa}H^{2}\widehat{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda_{K}}\right)$
$\leq\frac{180}{\kappa}(B+3)^{2}d^{2}H^{2}\log\left(1+\frac{K}{d\lambda_{K}}\right)\log\left(\frac{K}{\delta}\right)$ (22)
where the first inequality holds by $sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}(\theta)}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda_{k}d}\right),\forall\theta\in\Theta$ in Lemma 6 and the second inequality holds by the configuration of $\widehat{\beta}_{k}$.
Combining (20), (21), and (22), we have with probability at least $1-\delta$,
$\operatorname*{Reg}(K)$ $\leq\sqrt{2dK\log\left(1+\frac{K}{d\lambda_{K}}\right)}+\frac{180}{\kappa}(B+3)^{2}d^{2}H^{2}\log\left(1+\frac{K}{d\lambda_{K}}\right)\log\left(\frac{K}{\delta}\right)$
$\leq\widetilde{\mathcal{O}}\left(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2}\right).$
This finishes the proof. ∎
Appendix F Omitted Proofs for Section 5
F.1 Useful Lemma
Lemma 11.
For any $k\in[K]$, $h\in[H]$, define the second-order approximation of the loss function $\ell_{k,h}(\theta)$ at the estimator $\widetilde{\theta}_{k,h}$ as ~$\widetilde{\ell}_{k,h}(\theta)=\ell_{k,h}(\theta_{k,h})+\langle\nabla\ell_{k,h}(\widetilde{\theta}_{k,h}),\theta-\widetilde{\theta}_{k,h}\rangle+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{H_{k,h}(\widetilde{\theta}_{k,h})}^{2}$. Then, for the following update rule
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{\ell}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}}^{2},$
it holds that
$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}\leq$ $2\eta\left(\sum_{i=1}^{k}\ell_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}\ell_{i,h}(\widetilde{\theta}_{i+1,h})\right)+4\lambda B$
@@ F.1 Useful Lemma
Proof.
Based on the analysis of (implicit) OMD update (see Lemma 16), for any $i\in[K]$, we have
~$\big{\langle}\nabla\widetilde{\ell}_{i,h}(\widetilde{\theta}_{i+1,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}\leqslant\frac{1}{2\eta}\left(\|\widetilde{\theta}_{i,h}-\theta_{h}^{*}\|_{\mathcal{H}_{i,h}}^{2}-\|\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\|_{\mathcal{H}_{i,h}}^{2}-\|\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\|_{\mathcal{H}_{i,h}}^{2}\right)$
According to Lemma 4, we have
$\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\ell_{i,h}\left(\theta_{h}^{*}\right)\leqslant\big{\langle}\nabla\ell_{i,h}(\widetilde{\theta}_{i+1,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}-\frac{1}{\zeta}\left\|\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\right\|_{\nabla^{2}\ell_{i,h}(\widetilde{\theta}_{i+1,h})}^{2},$
where $\zeta=\log(K+1)+4$. Then, by combining the above two inequalities, we have
$\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\ell_{i,h}(\theta_{h}^{*})$ $\leqslant\langle\nabla\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\nabla\widetilde{\ell}_{i,h}(\widetilde{\theta}_{i+1,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\rangle$
$+\frac{1}{\zeta}\Big{(}\|\widetilde{\theta}_{i,h}-\theta_{h}^{*}\|_{\mathcal{H}_{i,h}}^{2}-\|\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\|_{\mathcal{H}_{i+1,h}}^{2}-\|\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\|_{\mathcal{H}_{i,h}}^{2}\Big{)}.$
We can further bound the first term of the right-hand side as:
$\big{\langle}\nabla\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\nabla\widetilde{\ell}_{i,h}(\widetilde{\theta}_{i+1,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}$
=\ = $\big{\langle}\nabla\ell_{i,h}(\widetilde{\theta}_{i+1,h})-\nabla\ell_{i,h}(\widetilde{\theta}_{i,h})-\nabla^{2}\ell_{i,h}(\widetilde{\theta}_{i,h})(\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}$
=\ = $\big{\langle}D^{3}\ell_{i,h}({\xi}_{i+1})[\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}](\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}),\widetilde{\theta}_{i+1,h}-\theta_{h}^{*}\big{\rangle}$
[…]
where the second equality holds by the mean value theorem, the first inequality holds by the self-concordant-like property of $\ell_{i,h}(\cdot)$ in Proposition 1, and the last inequality holds by $\widetilde{\theta}_{i+1,h}$ and $\theta_{h}^{*}$ belong to $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, and ${\nabla^{2}\ell_{i,h}({\xi}_{i+1})}\preceq I_{d}$.
Then, by taking the summation over $i$ and rearranging the terms, we obtain
$\big{\|}\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\big{\|}_{\mathcal{H}_{k+1,h}}^{2}\leqslant$ $\zeta\left(\sum_{s=1}^{k}\ell_{s,h}\left(\theta_{h}^{*}\right)-\sum_{s=1}^{k}\ell_{s,h}(\widetilde{\theta}_{i+1,h})\right)+\|\widetilde{\theta}_{1,h}-\theta_{h}^{*}\|_{\mathcal{H}_{1,h}}^{2}$
$+6\sqrt{2}B\zeta\sum_{s=1}^{k}\big{\|}\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\big{\|}_{2}^{2}-\sum_{s=1}^{k}\big{\|}\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\big{\|}_{\mathcal{H}_{i,h}}^{2}$
[…]
$+6\sqrt{2}B\zeta\sum_{s=1}^{k}\big{\|}\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\big{\|}_{2}^{2}-\sum_{s=1}^{k}\big{\|}\widetilde{\theta}_{i+1,h}-\widetilde{\theta}_{i,h}\big{\|}_{\mathcal{H}_{i,h}}^{2},$
where the last inequality holds by $\|\widetilde{\theta}_{1,h}-\theta_{h}^{*}\|_{\mathcal{H}_{1,h}}^{2}\leq\lambda\|\widetilde{\theta}_{1,h}-\theta_{h}^{*}\|_{2}^{2}\leq 4\lambda B$. Plugging $\zeta=2\eta$ finishes the proof. ∎
F.2 Proof of Lemma 2
Proof.
According to Lemma 11, we have
$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}^{2}\leq$ $2\eta\left(\sum_{i=1}^{k}\ell_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}\ell_{i,h}(\widetilde{\theta}_{i+1,h})\right)+4\lambda B$
We bound the right-hand side of the above lemma separately in the following. The most challenging part is to bound the term $\sum_{i=1}^{k}\ell_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}\ell_{i,h}(\widetilde{\theta}_{i+1,h})$. At first glance, this term appears straightforward to control, as it can be observed that $\theta_{h}^{*}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\bar{\ell}_{h}(\theta)\triangleq\mathbb{E}{y_{i,h}}[\ell_{i,h}(\theta)]$, where the instantaneous loss $\ell_{i,h}(\theta)$ serves as an empirical observation of ¯$\bar{\ell}_{h}(\theta)$. Consequently, the loss gap term seemingly can be bounded using appropriate concentration results. However, a caveat lies in the fact that the update of the estimator $\widetilde{\theta}_{i+1,h}$ depends on the information $\ell_{i,h}$, or more precisely $y_{i,h}$, making it difficult to directly apply such concentration results.
To address this issue, we decompose the loss gap into two components by introducing an intermediate term. Specifically, we define the softmax function as $[\sigma_{k,h}(z)]_{s}=\frac{\exp([z])_{s}}{1+\sum_{s\in\dot{\mathcal{S}}_{k,h}}\exp([z])_{s}},\forall s\in\mathcal{S}_{k,h}$. Using this definition, the loss function can be rewritten as:
@@ F.2 Proof of Lemma 2
$\ell_{k,h}(z_{k,h},y_{k,h})=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\operatorname{\mathds{1}}[y_{k,h}^{s^{\prime}}=1]\log\left(\frac{1}{[\sigma_{k,h}(z_{k,h})]_{s^{\prime}}}\right).$
@@
Define a pseudo-inverse function of $\sigma_{k,h}(\cdot)$ as $[\sigma_{k,h}^{-1}(p)]_{s^{\prime}}=\log\left(\frac{[p]_{s^{\prime}}}{1-\lVert p\rVert_{1}}\right),\forall p\in\{p\in[0,1]^{S_{k,h}}\mid\lVert p\rVert_{1}<1\}$. Then, the loss gap term can be decomposed into two parts as follows.
$\sum_{i=1}^{k}\left(\ell_{i,h}(\theta_{h}^{*})-\ell_{i,h}(\widetilde{\theta}_{i+1,h})\right)$
=\ = $\underbrace{\sum_{i=1}^{k}\left(\ell_{i,h}(\theta_{h}^{*})-\ell_{i,h}(z_{i,h},y_{i,h})\right)}_{\mathtt{term~{}(a)}}+\underbrace{\sum_{i=1}^{k}\left(\ell_{i,h}(z_{i,h},y_{i,h})-\ell_{i,h}(\widetilde{\theta}_{i+1,h})\right)}_{\mathtt{term~{}(b)}}$
@@
The design of the intermediate term was originally proposed by Zhang and Sugiyama [2023] in their study of the multiclass logistic bandit problem and was subsequently applied to the multinomial logit bandits problem by Lee and Oh [2024]. Notably, the intermediate loss is independent of the information contained in $y_{i,h}$, enabling the application of concentration results. Specifically, based on Lemma F.2 and Lemma F.3 of Lee and Oh [2024], we obtain the following upper bounds for them.
For term (a), let $\delta\in(0,1]$ and $\lambda\geq\max\{2,72cd\}$, for all $k\in[K],h\in[H]$, with probability at least $1-\delta$, we have
$\mathtt{term~{}(a)}$
@@
For term (b), for all $k\in[K],h\in[H]$, we have
$\mathtt{term~{}(b)}\leq\frac{1}{2c}\sum_{i=1}^{k}\left\|\widetilde{\theta}_{i,h}-\theta_{i+1,h}\right\|_{\mathcal{H}_{i,h}}^{2}+\sqrt{6}cd\log\left(1+\frac{k+1}{2\lambda}\right).$
Combining term (a) and term (b), we have